OpenStudy (anonymous):

Could someone explain to me how to solve this: Simplify: (64x^5)/(4x^2) ^ is "to the power of" Thanx!!!

OpenStudy (anonymous):

Divide the numbers, subtract the exponents. 64 / 4 = 16 and x^5 / x^2 = x^3 Then put it together. Final answer : 16x^3

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

Any time!